OpenStudy (zenmo):

Verify the Trigonometry Identity. (cos4x+cos2x) / (sin4x+sin2x) = cot3x. Much Appreciated! @@QUESTION ON HOLD@@

OpenStudy (lxelle):

Do you know how to apply the double angle formula?

OpenStudy (zenmo):

Kind of...

OpenStudy (lxelle):

Show me your attempt.

OpenStudy (anonymous):

@LXelle Is there a trig identity for cos4x? O_O

OpenStudy (lxelle):

Did I mention double angle?

OpenStudy (zenmo):

Yes

OpenStudy (zenmo):

\[\cos(4x)=\cos ^{2}(x)-\sin ^{2}(2x)=2\cos ^{2}(2x)-1=1-2\sin ^{2}(2x) \]

OpenStudy (zenmo):

\[\sin(4x)=2\sin(2x)\cos(2x)\]

OpenStudy (zenmo):

\[1-2\sin ^{2}(2x)+1-2\sin ^{2}(x) / 2\sin(2x)\cos(2x)+2\sin(x)\cos(x) ??\]

OpenStudy (lxelle):

How about using product rule for the numerator and denominator?

OpenStudy (zenmo):

like sinUsinV=1 / 2 [ cos (u-v) - cos (u + v) ]?

OpenStudy (lxelle):

yeah yeah something like that.

OpenStudy (zenmo):

Going to try it then, putting this question on hold again while I work on another attempt.

OpenStudy (lxelle):

In your case the numerator would be 2 cos ((4x+2x)/2) cos ((4x-2x)/2)